OpenStudy (anonymous):

Consider the line y=-4x+2. Find the equation of the line that is perpendicular to this line and passes through the point (6,-6) . Find the equation of the line that is parallel to this line and passes through the point (6,-6) .

OpenStudy (anonymous):

-6=1/4(6)+b -6=3/2+b -15/2=b y=1/4x-15/2 (perpendicular)

OpenStudy (anonymous):

-6=-4(6)+b -6=-24+b 18=b y=-4x+18 (parallel)

hartnn (hartnn):

thats correct

OpenStudy (anonymous):

woot thanks @hartnn

hartnn (hartnn):

both correct, good! welcome ^_^